• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 07:55:41 on July 11, 2010

    Entry number 331433

    keyword=Shift Summary (OWL)


    Shift Summary for July 11, 0000-0800


    Shift Crew:

    Shift Leader: Brad Sawatzky
    Target Operator: Mahbub Khandaker

    Kinematics (nominal setpoints):

    Beam Energy: 2.2619 GeV
    Beam Current: 120 uA
    Left HRS: 1.1310 GeV/c
    Right HRS: 1.1310 GeV/c
    Septum Current: 775.6 A

    Summary:


    Time    Description
    0000    Started shift with Right Q1 down and unable to reset.  Heidi came
            in to look at the problem.  RCs for Hall A and C decided to
            opportunistically shift Hall C's 3hr test plan from 6am to start now
            (and end at 3am) instead.  There was also an issue with the RF
    seperator,
            but that seems to have resolved by 00:15.
    0130    Right Q1 is back up.  Heidi swapped in the previous control board so
            we can turn the regulation back on.
    0305    Bringing beam back to Hall A.  We're back in production.
    0600    Stable beam up to this point.  Got a little trippy starting at 6am.
    0640    MCC called. They accidentally cycled our dipole string after the last
            trip (hit the wrong button).  We're back up after a few minutes.
    0650    Beam stability seems to have improved.
    0735    Beam twitched a couple of mm in BPMB x and y (very short transient).
            Rad44 monitors jumped a bit, but no FSD trip.  Beam stable before and
            after the blip.
    

    Run List:


    Run     Target      Events      Comments
    2037      Ta                    Restarting production, spot++ run
    (different prescales)
    2038      Ta        500k        Production settings
    2039      Ta        500k        Production settings
    2039      Ta        500k        Production settings
    2040      Ta        500k        Production settings
    2041      Ta        500k        Production settings
    2042      Ta        174k        Production settings, stopped early due to
    extended trip
    2043      Ta        500k        Production settings       
    2044      Ta        In progress at 0800... (